OpenStudy (anonymous):

-3x^2+4y^2-(5x^2-7y^2-1) now i came up with the answer of 32

OpenStudy (angela210793):

-3x^2+4y^2-(5x^2-7y^2-1)=-3^2+4y^2-5x^2+7y^2+1=-8x^2+11y^2+1

OpenStudy (angela210793):

How did you get 32? O.o

OpenStudy (anonymous):

-3x^2+5x^2 (4y^2-7y^2

OpenStudy (angela210793):

no u don't multiply them...u just add and subtract those just as with natural no.s

OpenStudy (angela210793):

Bella is it minus before the bracket?>

OpenStudy (anonymous):

yes thats what it looks like

OpenStudy (anonymous):

\[11y^2-8x^2+1\]

OpenStudy (angela210793):

then u don't have to multiply anything -3-5=-8 right? (integers) u don't have these u have -3x^2=5x^2=-8x^2

OpenStudy (anonymous):

yes its definitaly minus

OpenStudy (angela210793):

so it's =-8x^2+11y^2+1

OpenStudy (anonymous):

now these are to the 2nd power i couldn't figure out how to type it in

OpenStudy (angela210793):

write it by using the blue button below ''equation''

OpenStudy (anonymous):

Yes it's \[11y^2-8x^2+1\] because \[(-3x^2-5x^2)(+4y^2+7y^2)+1\]Then rearrange so the highest coefficient is to the left giving the above

OpenStudy (anonymous):

The trick is to change the sign of all the terms in the bracket when removing them thus the terms in the bracket become:\[-5x^2+7y^2+1\]

OpenStudy (anonymous):

thanks angela i never thought about using that

OpenStudy (angela210793):

Lol :P:P

OpenStudy (anonymous):

okay Kevsturge this makes sense

OpenStudy (anonymous):

so the answer is 25 then and not 32

OpenStudy (anonymous):

how did you get 25?

OpenStudy (angela210793):

Bella can u pls write the problem as it is written in ur book? this way we may help u better :)

OpenStudy (anonymous):

okay i am lost am adding here or subtracting

OpenStudy (anonymous):

okay one minute

OpenStudy (anonymous):

Ah, I am missing something lol

OpenStudy (anonymous):

\[-3^2+4y ^2-(5^2-7y ^{2}-1)\]

OpenStudy (anonymous):

3x sory

OpenStudy (anonymous):

and 5x

OpenStudy (anonymous):

Sure does this equal 0 or something otherwise I cannot get an answer to this other than to collect the terms.

OpenStudy (anonymous):

i am starting to wonder the same thing

OpenStudy (anonymous):

Does the question say solve for x or y or something else?

OpenStudy (angela210793):

so then it's the same answer we gave u if it is...\[-3x^2+4y^2-(5x^2-7y^2-1)=-3x^2+4y^2-5x^2+7y^2+1=-8x^2+11y^2+1\]

OpenStudy (anonymous):

sorry got to go, will be back later..

OpenStudy (angela210793):

bye :)

OpenStudy (anonymous):

omg its after 2pm i have been so caught up thks guys i have to get ready for work i will be back later man

OpenStudy (anonymous):

tks Angela i couldn't have figured this one out

OpenStudy (angela210793):

You're welcome :):) Have a gr8 day :):)

OpenStudy (anonymous):

will you be on later tonight

OpenStudy (angela210793):

I don't think so...I have to study conics :(:(:( I'm very behind :(:( Sorry....other ppl will help you :):)

OpenStudy (anonymous):

okay tks have a bless one

OpenStudy (angela210793):

u 2 :)